OpenStudy (anonymous):

If a cone has a height of 5 inches and a volume of 135pi, in^3find the diameter of the circular base.

OpenStudy (anonymous):

Volume is cone=1/3x\[\pi\]xr^2xh 135\[\pi\]= 1/3x\[\pi\]x r^2x 5 (135x3)/5= r^2 27x3=r^2 81=r^2 \[\sqrt{81}\] = r 9=r , (diameter is 81, radius is 9)